INAF queue observing, 2015 Mar 22-23

Observer: R. Speziali, S. Bisogni
Telescope Operator: D. Gonzales Huerta
Support Astronomer: Olga on call
Instrument: LUCI - MODS -LBC-R

Summary

At the sunset the sky is mostly clear, humidity 67%.

Night duration (12\xB0 el twilight) 19:30-05:30 = 11 hrs

Observing time: 11 hrs

Weather time loss: 0 hrs (0%)

Data Summary

At the beginning of the night the sky is mostly clear

Project_RUN Objects Filter ExpTime Notes
Piconcelli B1 telluric HIP52877 grating 210_zJHK, filter H 4s  
Piconcelli B1 1106+6400 grating 210_zJHK, filter H 30min COMPLETED
Piconcelli B2 telluric HIP58396 grating 210_zJHK, filter H 4s  
Piconcelli B2 1157+2724 grating 210_zJHK, H 20min  
Fiore B goodsnorth 2 dual grating + i-sloan (LBC-R) 1.30 h 30 min out of specs for diff refraction?
Fiore B standard Feige34 dual grating 3min  
Fiore A goodsnorth 1 dual grating + i-sloan (LBC-R) 1h COMPLETED
Mignoli MOS21 SDSSJ1411 M21 red grating 3h  
Mignoli LBC MOS20_r J1411+12 filter z-sloan 30m  
Mignoli LBC MOS20_r J1411+12 filter i-sloan 30m  
Mignoli LBC MOS20_r J1411+12 filter r-sloan 15m  
Luci Calibrations        

Details

We run LUCI darks: executeLUCIscript.sh DoDarks.txt luci.20150323.0002-0015

01:44 the TO opens the dome

slewing to Piconcelli B1

02:05 executeLUCIscript tellB1_LS_JHeI.acq luci.20150323.0016-0022

02:24 executeLUCIscript tellB1_LS_210_H.spec luci.20150323.0023-0024

02:27 executeLUCIscript B1_LS_H.acq luci.20150323.0025-0028-

02:49 executeLUCIscript B1_LS_H.spec luci.20150323.0029-0034
  • seeing 0.7''

slewing to Piconcelli B2

03:23 executeLUCIscript.sh tellB2_LS_JHeI.acq luci.20150323.0035-0040

03:30 executeLUCIscript.sh tellB2_LS_JHeI.spec luci.20150323.0041-0042

03:35 executeLUCIscript.sh B2_LS_H.acq luci.20150323.0043-0046

03:55 executeLUCIscript.sh B2_LS_H.spec luci.20150323.0047-0048
  • seeing 0.9''

Switch to MODS - LBC_R

FIORE B 04:38 acqMODS goodsnorth2_mos.acq mods1r.20150323.0002-0004

04:47 running DOFPIA

04:50 execMODS goodsnorth2_mos_3.obs mods1r.20150323.0005-0008 mods1b.20150323.0002-0005
  • seeing 0.9'' at the beginning of the 3rd integration
  • 05:06* running DOFPIA again, it couldn't converge
  • it eventually converged
  • 05:11 launched goodsnorth_isloan_new
  • seeing 0.9'' at the beginning of the 4th integration

06:00 execMODS execMODS goodsnorth2_mos_4.obs mods1r.20150323.0009-0010 mods1b.20150323.0006-0007

06:40 acqMODS Feige34 mods1r.20150323.0011-0013

06:57 execMODS feige34_ls.obs mods1r.20150323.14 mods1b.20150323.0008

FIORE A 07:04 acqMODS goodsnorth1_mos.obs mods1r.20150323.15-17

07:18 execMODS goodsnorth1_mos_4.obs mods1r.20150323.0018-21 mods1b.20150323.0009-12
  • seeing 0.9'' at the beginning of the first integration
  • 07:33 running DOFPIA
  • seeing 0.9'' at the beginning of the 3rd integration
  • 08:11 running DOFPIA
  • seeing 1'' at the beginning of the last integration

Slewing to Mignoli MOS21

08:32 running DOFPIA

08:32 acqMODS mods.21..acq mods1r.20150323.0022-0025

08:50 execMODS mods.21.1.obs mods1r.20150323.0026-0029
  • seeing 1.1'' at the beginning of the first integration
  • 08:52 launched 1_MOS21_ z.ob
  • 08:59 running DOFPIA
  • seeing 0.9''at the beginning of the 2nd integration
  • seeing 0.9'' at the beginning of the 3rd integration
  • 09:38 LBC GUI got stuck; we ask the TO to restart the software
  • seeing 0.9'' at the beginning of the 4th integration

09:59 execMODS mods.21.2.obs mods1r.20150323.0030-0033
  • seeing 0.9 at the beginning of the first integration
  • 10:01 running DOFPIA; the TO has restarted the LBC software, so we have lost collimation
  • seeing 1.0'' at the beginning of the 2nd integration
  • 10:27 launched 2_MOS21_i.ob
  • seeing 1'' during the third integration
  • seeing 1'' at the beginning of the last integration

11:10 execMODS mods.21.3.obs mods1r.20150323.0034-0037
  • seeing 1''
  • 11:15 running DOFPIA
  • 11:19 launched 3_MOS21_i.ob
  • 11:22 cirri passing, the seeing is getting worse during the first integration (1.4'')
  • 11:44 launched 3_MOS21_i.ob (seeing = 1.2")
  • 12:09 launched 4_MOS21_z.ob

12:18 execMODS mods.21.4.obs mods1r.20150323.0038 only one spectrum AND TO BE REPEATED!!!
  • seeing 1'' at the beginning of the integration
  • both the last MODS and LBC integration are to be repeated!!! They are pretty much the same as flats! Sun is rising!
12:18 Luci Calibration
  • Piconcelli B1 FLAT (51-61) too many counts (30.000) so we set the exposure time to 7 sec and repeat the calibration
  • Piconcelli B1 FLAT (62-71)
  • Piconcelli A4 FLAT (74-83) these are flat with good counts (10.000-20.000)
  • Piconcelli A4 ARC (84-87)
  • Piconcelli B1-B2 ARC (88-91)

-- %USERSIG{SusannaBisogni - 2015-03-22}%

Comments

 
Topic revision: r7 - 29 Apr 2015, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback